# Dark Energy from Time Crystals

### v1

@@ 1. The time crystal region: $X\leq\dfrac{\kappa_{2}}{6\lambda_{2}}$
2. The forbidden region bounded below by the orbit of the time crystal $X_{t}$ and above by the zero group velocity $X_{d}$ respectively: $\dfrac{\kappa_{2}}{6\lambda_{2}}<X<\dfrac{\kappa_{2}}{2\lambda_{2}}$
  • $N<0$,
  • $D>0$,
[…]
@@ 3. The quintessence region: $X\geq X_{d}=\dfrac{\kappa_{2}}{2\lambda_{2}}$
There is an additional point of interest in the forbidden region $\it{2}$, where the speed of sound goes through zero again and halfway in this region it switches from negative to positive indicating stability, at $X=X_{0}=\dfrac{\kappa_{2}}{4\lambda_{2}}=X_{d}/2$ found by the condition of Section $\it{1}$: $2Xg"(X)\leq-g^{\prime}(X)$. In the forbidden region $X_{t}\leq X\leq X_{0}$ the momentum $\Pi$ obtained by the quadratic equation above $\rho\geq 0$ is imaginary, again indicating that the field would be damped to zero quickly in this region. Therefore, stability ($c_{s}^{2}\geq 0$) occurs only when the field is confined in the time crystal region or when it is specially prepared to start at high conjugate moment for $X\geq X_{d}$:
@@
However, the unstable region above $X_{t}$ is inaccessible to the field since time crystals are in stable bound states and the speed of sound squared diverges at the boundary $X=X_{t}$, as given by cusps in the momentum space. These imply that the field is confined in the crystal region and it would require infinite work done to it to jump over the crystal boundary and to break the crystal structure. Even if the field were to tunnel through $X_{t}$ boundary with a small probability it would die out to zero quickly since the momentum becomes imaginary in the forbidden region. In other words, these crystals are very stable structures.
@@
$w_{\phi}<-1$ (phantom behavior), $c_{s}^{2}\geq 0$ (stability), $\rho\geq 0$ if the potential $V(\phi)$ is sufficiently large to ensure $\rho\geq 0$.The latter is automatically satisfied throughout the field space if we require $W(\phi)\geq 0$, which can be easily achieved by multiplying the potential $V(\phi)$ by an overall constant factor without changing or tuning its functional form.
For $X\geq\dfrac{\kappa_{2}}{2\lambda_{2}}$, stability is maintained and $\rho\geq 0$, but $w_{\phi}>-1$, so the phantom behavior is lost and replace by a quintessence type dynamics. However, if the field is bound below $X_{d}$ in the time crystal region it would not be possible for the field to break away from its bound state and suddenly jump over to $X=X_{d}$ which is the boundary of the quintessence region since the momentum $\Pi_{d}=0$ and the group velocity are zero at that initial boundary. The field could not access the region $X>X_{t}$.
Let’s classify the points and regions of interest in field space in the table below.
Based on the equation of motions for $X$, the behavior of $w_{\phi}$, $c^{2}_{s}$, and $\rho$ we showed that the field space is divided into distinct regions of interest, one of which contains the time crystals. Table 1 summarizes these results, detailing the range of $X$, energy density, equation-of-state parameter, and sound speed squared $c_{s}^{2}$ for all the field space regions.
@@
This indicates stable perturbations within this interval. At the boundary $X=X_{t}$ the energy density reaches:
@@
Here, the equation-of-state parameter $w_{\phi}$ attains its minimum value within the phantom regime ($w_{\phi}<-1$). The sound speed squared $c_{s}^{2}$ approaches infinity at the crystal’s boundary, indicating a stable time crystal structure.
The field would require an infinitely large pressure gradient, or force, to break out of the crystal boundary $X_{t}$. In a sense, this boundary acts as a hard wall potential for the forbidden region $X>X_{t}$, with the field trapped on its other side $X\leq X_{t}$. Hence, the region $X_{t}<X<X_{d}=\frac{\kappa_{2}}{2\lambda_{2}}$ is inaccessible to the field. The speed of sound squared being greater than one is typical of crystalline structures. Its contribution to the metric perturbation equation is to dampens perturbations that would contribute to cold dark matter, thereby offering a smoking gun for testing this model.
@@
indicating stable perturbations within the quintessence regime. However, for the reasons discussed above, regions above the boundary of the time crystal with $X\ >X_{t}$ are not accessible unless the field is especially prepared to start at high momentum. Even then, the time crystal region $X\leq X_{t}$ is favored as the state that minimize energy.
@@ IV Conclusions
In this work, we investigated the time crystal model proposed in  [6, 7] and showed that in the context of cosmology, it gives rise to a phantom dark energy ideal fluid, with an equation of state $w_{\phi}\leq-1$, a positive speed of sound squared $c_{s}^{2}\geq 0$, and positive energy density $\rho\geq 0$, which drives the expansion of the universe towards the Big Rip.
Furthermore, the speed of sound in the crystal region $X\leq X_{t}$ is such that $c_{s}^{2}\geq 1$ typical of crystalline structures. Observationally this sound speed would serve to dampen cold dark matter perturbation providing a way of testing this model.
The speed of sound squared diverges at the boundary of the time crystal in field space $X_{t}$ which implies that these are very stable structures bound by a hard wall in field space, such that the field would require vast amount of work to be done for it to break away from the crystal region. However, if for some reason, the field was prepared to start at high conjugate momentum or it tunneled through the hard wall, it would enter a forbidden region $X_{t}<X<X_{d}$ where the field is exponentially decaying, separating it from the quintessence region $X>X_{d}$. Yet, not only is the field confined to remain in the crystal region $X\leq X_{t}$ but dynamics favors it to naturally be found there since the crystal state minimizes energy.
@@
Acknowledgment: L. Mersini-Houghton is grateful to Y. J. Ng, O. Akarsy, E. Di Valentino for useful discussions, and to the Klingsberg foundation and the Bahnson fund for their support.

### v2

@@ 1. The time crystal region: $X\leq\dfrac{\kappa_{2}}{6\lambda_{2}}$
2. The forbidden region bounded below by the orbit of the time crystal $X_{t}$ and above by the locus of the Lagrangian $X_{d}$ respectively: $\dfrac{\kappa_{2}}{6\lambda_{2}}<X<\dfrac{\kappa_{2}}{2\lambda_{2}}$
  • $p_{X}<0$,
  • $\rho_{X}>0$,
  • $c_{s}^{2}<0,$ (unstable) goes through zero at $X=X_{0}$, and $\rho>0,w_{\phi}<-1$.
@@ 3. The quintessence region: $X\geq X_{d}=\dfrac{\kappa_{2}}{2\lambda_{2}}$
There is an additional point of interest in the forbidden region $\it{2}$, where the speed of sound goes through zero again and halfway in this region it switches from negative to positive indicating stability, at $X=X_{0}=\dfrac{\kappa_{2}}{4\lambda_{2}}=X_{d}/2$ found by the condition of Section $\it{1}$: $2Xg"(X)\leq-g^{\prime}(X)$. In the forbidden region $X_{t}\leq X\leq X_{0}$ the momentum $\Pi$ obtained by the quadratic equation above $\rho\geq 0$ is imaginary and the field solution is a decaying mode. Therefore, stability ($c_{s}^{2}\geq 0$) occurs only when the field is confined in the time crystal region or when if it were to start at high conjugate moment for $X\geq X_{d}$:
@@
However, the unstable region above $X_{t}$ is inaccessible to the field since time crystals are in stable bound states and the speed of sound squared diverges at the boundary $X=X_{t}$, as given by cusps in the momentum space. These imply that the field is confined in the crystal region and it would require infinite work done to it to jump over the crystal boundary and to break the crystal structure. Even if the field were to tunnel through $X_{t}$ boundary with a small probability it would die out to zero quickly since the momentum becomes imaginary and the field solution is a decaying mode in the forbidden region. In other words, these crystals are very stable structures, and the field is confined within the boundaries of the crystal.
@@
$w_{\phi}<-1$ (phantom behavior), $c_{s}^{2}\geq 0$ (stability), $\rho\geq 0$ if the potential $V(\phi)$ is sufficiently large to ensure $\rho\geq 0$.The latter is automatically satisfied throughout the field space if we require $W(\phi)\geq 0$, which can be easily achieved by multiplying the potential $V(\phi)$ by an overall constant factor without changing or tuning its functional form. Despite that $p_{X}$ and $\rho_{X}$ are negative within the crystal boundaries, the important factor for the stability of the field is their ratio which gives the speed of sound squared be positive, $c_{s}^{2}=\frac{p_{X}}{\rho_{X}}>0$. As can be seen from the field equation in an expanding FRW universe, Eqn. (35), the Hubble drag term multiplied by the speed of sound overwhelmingly dominates over the potential term. The only effect of $\rho_{X}<0$ is to invert the sign of the potential, a term which is insignificant relative to the growing Hubble drag term, ( since $H\simeq a(t)^{\frac{-3}{2(1+w)}},c_{s}^{2}\geq 1$. (This stability remains true for the field perturbations also, as studied in  [10].)
For $X\geq\dfrac{\kappa_{2}}{2\lambda_{2}}$, stability is maintained and $\rho\geq 0$, but $w_{\phi}>-1$, so the phantom behavior is lost and replace by a quintessence type dynamics. However, if the field is bound below $X_{d}$ in the time crystal region it would not be possible for the field to break away from its bound state and suddenly jump over to $X=X_{d}$ which is the boundary of the quintessence region, with the momentum $\Pi_{d}=0$ and the group velocity are zero at that initial boundary. The field could not access the region $X>X_{t}$.
The table below 1 classify the three regions of interest in field space, and summarizes these results, detailing the range of $X$, energy density, equation-of-state parameter, and sound speed squared $c_{s}^{2}$ for all the field space regions, as discussed above.
@@
This ensures stable perturbations within this interval. At the boundary $X=X_{t}$ the energy density reaches:
@@
Here, the equation-of-state parameter $w_{\phi}$ attains its minimum value within the phantom regime ($w_{\phi}<-1$). The sound speed squared $c_{s}^{2}\rightarrow\infty$ at the crystal’s boundary, indicating a stable time crystal structure which confines the field within.
The field would require an infinitely large pressure gradient, or force, to break out of the crystal boundary $X_{t}$. In a sense, this boundary acts as a hard wall potential for the forbidden region $X>X_{t}$, with the field trapped on its other side $X\leq X_{t}$. Hence, the region $X_{t}<X<X_{d}=\frac{\kappa_{2}}{2\lambda_{2}}$ is inaccessible to the field.
The speed of sound squared being greater than one within the crystal region, is typical of crystalline structures. Its contribution to the metric perturbation equation is to dampens perturbations  [9] that would contribute to cold dark matter, thereby offering a smoking gun for testing this model.
@@
indicating stable perturbations within the quintessence regime. However, for the reasons discussed above, regions above the boundary of the time crystal with $X\ >X_{t}$ are not accessible due to the hard wall at the boundary $X_{t}$ of the crystal where $c_{s}^{2}\rightarrow\infty$. Perturbations around the crystal solution, ($\ddot{\delta\phi}+\frac{c_{s}^{2}k^{2}}{a^{2}}+(3H+\frac{6\lambda_{2}\ddot{\phi}\dot{\phi}}{\rho_{X}})\dot{\delta\phi}+\frac{1}{\rho_{X}}\frac{\partial^{2}V}{\partial\phi^{2}}=0$), where studied in [10] and shown to be stable. Besides, the time crystal region $X\leq X_{t}$ is favored as the state that minimize energy.
@@ IV Conclusions
Until recently, phantom dark energy models have been studied in the context of k-essence fluids with non-canonical kinetic energies. In this work, we investigated the time crystal model proposed in  [6, 7] and showed that in the context of cosmology, it gives rise to phantom dark energy, with an equation of state $w_{\phi}\leq-1$, a positive speed of sound squared $c_{s}^{2}\geq 0$, and positive energy density $\rho\geq 0$, which drives the expansion of the universe towards the Big Rip. While both, k-essence models and the time crystal model discussed here, have in common non-canonical kinetic terms, in the latter case we are dealing with a collection of crystalline structures that break time symmetry and behave as a phantom fluid instead of matter. Furthermore, in contrast to k-essence models where the field explores a large volume in field space, in time crystals the field is confined within the boundary of the crystal whereby the rest of the field space is inaccessible by a hard wall.
The possibility of ghost or gradient instabilities are often a cause of concern in k-essence models. These instabilities on the field and its perturbations do not arise in the present model. The time crystals are on an expanding FRW background where the ever increasing Hubble drag term, multiplied by a crystal’s positive speed of sound squared, dominates over the potential term.
Furthermore, the speed of sound in the crystal region $X\leq X_{t}$ is such that $c_{s}^{2}\geq 1$ typical of crystalline structures. Observationally this sound speed would modify the Newtonian potential in a unique way, which could provide a way of testing this model. The couple equations of metric perturbations and field equations can only be solved numerically and we will report this results in an upcoming paper.
The speed of sound squared diverges at the boundary of the time crystal in field space $X_{t}$ which implies that these, and their perturbations, are very stable structures bound by a hard wall in field space. Yet, not only is the field confined to remain in the crystal region $X\leq X_{t}$ but dynamics favors it to naturally be found there since the crystal state minimizes energy.
@@
Acknowledgment: L. Mersini-Houghton is grateful to Y. J. Ng, O. Akarsu, E. Di Valentino for useful discussions, and to the Klingsberg foundation and the Bahnson fund for their support.